OpenStudy (anonymous):

-0.6x^2-6x+30 = 0 solve by completing the square?

OpenStudy (anonymous):

cannot solve by factoring. Must use the formula to get the answer

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

u're welcome

OpenStudy (anonymous):

Can it be solved by completing the square?

OpenStudy (anonymous):

Yes it can. wait a minute

OpenStudy (anonymous):

-0.6x^2 - 6x + 30 = 0 -0.6(x^2 + 10x - 50) = 0 x^2 + 10x - 50 = 0 (x^2 + 10x + ___) = 50 (x^2 + 10x + 25) = 50 + 25 (x + 5)^2 = 75 x + 5 = + or - sqrt(75) x = -5 + 5sqrt(3) or x = -5 - 5sqrt(3)